# Evolution of Photonic Quantum Machine Learning under Noise

### 10 Mar 2026 v1

@@ 5.2 Advanced Characterisation Techniques
After characterizing the noise using above techniques, we need to consider developing effective noise mitigation strategies to reduce the impact of noise on photonic quantum machine learning systems.
@@ 6.1 Hardware-Level Mitigation
Mitigation Strategies at the Hardware Level focus on improving the quality of the physical components that comprise the Photonic Quantum System as a method for eliminating Noise. One of the most effective methods to reduce Noise is by incorporating quality Optical Components like Optical Waveguides with High-efficiency Single Photon Detectors, and Stable Laser Sources with Low Losses. By implementing Low Loss Waveguides, High-efficiency Single Photon Detectors, and Stable Laser Sources, it becomes possible to effectively reduce the number of photons detected due to inefficiencies at the Detector and lost to the environment due to extinction of the light [45]. Likewise, the incorporation of Active Stability Techniques like Feedback Control Systems will further enhance the stability of Optical Components thereby eliminating Phase Noise [25].
@@ 6.2 Encoding-Based Mitigation
Encoding-Based Mitigation techniques are focused on creating a set of specific types of encoded agencies to help protect quantum information from degrading environmental disturbance or noise. One encoding method for Quantum Information is called Redundant Encoding, which is a way of spreading Quantum Information throughout multiple Photons or Modes of light within a physical Composite quantum [11]. Redundant structured encoding spreads the Quantum Information out, providing more flexibility with regards to loss of Photons. The second form of Encoding method is referred to as the use of Decoherence-Free Subspaces(DFS), which involves utilizing specific subspaces in the Quantum Region in order to encode the Quantum Information with immunity from a number of different forms of environmental noise, specifically forms of noise created through the combined dephasing of our universe [31]. Another encoding technique for encoding Quantum Information is the implementation of Quantum Error-Correcting Codes, which detect and correct errors that occur through environmental noise during Quantum Processes. Each of the above Encoding techniques work to increase the resilience of Quantum Machine Learning Systems to a combination of different environmental noise sources.
@@ 6.3 Algorithmic Mitigation
Algorithmic Mitigation techniques aim to build Quantum Machine Learning Algorithms that can stay resiliently when faced with noise in Photonic Quantum Systems. To that end, Error Mitigation Methods are available; these are designed to help eliminate the effects of noise on the output of quantum algorithms and therefore do not have to perform complete error correction. Error Mitigation Methods include: Probabilistic Error Cancellation, Zero-Noise Extrapolation, and Symmetry Verification [47]. Another approach used by Algorithmic Mitigation is to create Noise-Resilient Algorithms that have been created specifically to handle certain noise types well. Noise-aware training is a technique that allows a user to train an algorithm to understand what type of noise is present during its learning phase. By combining these Algorithmic Mitigation Techniques together, we are able to improve Quantum Machine Learning Algorithm performance despite the presence of noise. Other than above approches, Machine Learning-Based Noise Reduction techniques can also be used to reduce noise in photonic quantum systems. These techniques use machine learning models to identify and filter out noise from quantum data, thereby improving the quality of the quantum information being processed [49]. “Figure. 8” illustrates a such machine-learning–based quantum error mitigation approach.
Figure 8: Machine-learning–based quantum error mitigation, where a trained model maps noisy expectation values from a quantum circuit to noise-free predictions without requiring additional mitigation circuits [30].
@@ 6.4 Hybrid Classical-Quantum Approaches
Hybrid Classical-Quantum Approaches are strategies that merge Classical and Quantum Computing technologies to reduce this interference. A Hybrid Classical-Quantum Approach can also be achieved through Classical Post-Processing, which uses Classical Algorithms for analysis and correction of the output data from Quantum Algorithms that have been degraded by noise as a result of their execution in a noisy environment [47]. On the other hand, Variational Hybrid Algorithms can also be a Hybrid Algorithm based on a combination of Quantum Computing and Classical Optimization Techniques to develop adaptive algorithms that can continually adjust to the noise throughout their execution [46]. Hybrid Classical-Quantum Approaches use a combination of the advantages of both Classical and Quantum Computing to develop effective ways to reduce the effect of noise in the development of Photonic Quantum Machine Learning Systems.
@@ 7 Recent Advances and Experimental Demonstrations
2016 was a seminal year for continuous-variable (CV) photonic quantum computing, as it was the year in which a 1 million-mode photonic cluster state was demonstrated, thus demonstrating the scale-up capabilities of CV photonic systems [51]. Using this very large-scale cluster state, a variety of quantum information processing tasks were also demonstrated (e.g., quantum teleportation, entanglement distribution).
Starting in 2021, research teams have developed integrated programmable photonic processors at multiple institutions (Xanadu, NIST, MIT), thus enabling multiple quantum algorithms to be implemented on a single photonic chip, thus enabling reconfigurable and scalable photonic quantum computing [8]. In addition, the integrated programmable photonic processors also reduced system noise by decreasing both the number of optical components and improving the stability of the overall system. “Figure. 9” illustrates a overview of a processor.
Figure 9: Overview of an integrated programmable photonic processor [8].
@@
To continue developments toward 2024, a new superconducting-photonics hybrid chip was conceived, allowing for  99% loss resilience at photon levels, and leading to the reduction of the leading type of noise in photonic systems [28]. These developments will allow for a much larger scale of practical deployment of photonic-based quantum computing.
@@ 8 Discussion
According to above sections, we can discuss about three main points such as theoretical frameworks and noise modeling perspectives, near-term applications of PQML under noise, and long-term prospects and open challenges.
@@ 8.1 Theoretical Frameworks and Noise Modeling Perspectives
To develop the most promising strategies to reduce or eliminate noise in photonic quantum systems, we must construct a theoretical foundation that accurately captures the types of noise found in these systems. Noise may have many different effects on the state of a quantum system; thus, there are several types of noise models that have been proposed to define how it impacts a quantum state, such as amplitude and phase damping, and depolarizing channels [13]. By accurately modelling the various noise processes affecting photonic quantum systems, we may be able to predict how each noise type will act upon different types of quantum algorithms, as well as design robust quantum algorithms that are resilient to all types of noise. Furthermore, advanced techniques for modelling noise, including the development of stochastic noise models and non-Markovian noise models [12], provide the most accurate representation of the many types of noise that exist in the real world and will allow us to study the temporal correlations in noise, the memory effects associated with it, and the long-term behaviour of quantum states in noisy environments.
@@ 8.2 Near-Term Applications of PQML under Noise
Advances in photonic quantum machine learning (PQML) will bring about a transformational impact in many areas including but not limited to areas that are susceptible to noise. For example due to its nature of being able to classify chemicals, photos, videos, etc. in real time, PQML could allow for more precise image classification than classical classification techniques [34]. A good example of this would be using a PQML algorithm for classifying chemicals found in samples of murky water which cannot be done effectively by current methods of classification. As PQML continues to evolve researchers will find even greater opportunities to utilize the incredible capabilities of PQML to their advantage; particularly due to the inherent robustness of PQML digital circuitry and room temperature operation.
@@ 8.3 Long-Term Prospects and Open Challenges
The future of photonic quantum machine learning (PQML) looks bright, but many Current challenges are preventing this Field from reaching its maximum potential. Among these challenges are the limitations on the scalability of quantum photon systems. Although considerable work has been completed in building integrated photonic circuits, creating systems that may undertake more qubts or continue functioning at a low noise level when built larger will be difficult to develop [8]. Additionally, integrating photonic quantum systems with other methods for executing quantum technologies such as superconducting qubit systems or an ion confined to a magnetic field presents not only opportunities to use the advantages of each unique quantum scheme, but also creates additional issues associated with noise control and system coherence [7]. Therefore, addressing these challenges will require increased interdisciplinary cooperation among physicist engineers and computer scientists to advance the state-of-the-art of PQML further.
@@ 9 Conclusion
Our paper provided a comprehensive review of the state of research into Quantum Machine Learning (QML) and Photonic Quantum Computing (PQC) that focused on the noise limitation in the quantum machine learning systems. We have covered the general principles related to the photonic quantum machine learning, including the various types of Photonic Quantum Computing Architectures, algorithms developed using Quantum Machine Learning on Photonic platforms, various noise sources in Photonic Quantum systems and their classifications, including natural noise mechanisms and the characteristics associated with the Photonic Quantum Systems. Also, we provided an overview of how the various forms of noise affect the quantum machine learning algorithm/algorithms and discussed the current research into Algorithm-Specific Noise Sensitivity and the design of noise-resilient algorithms.
Additionally, this review has included an overview of the various types of methods and current status of research into the various methods and current research into Noise Characterization, as well as how the various techniques, including hardware-level techniques to develop encoding-based, algorithmic and hybrid classical-quantum methods, and experimental demonstrations. Overall, this article highlights not only the challenges and achievements in managing the problems caused by noise within photonic quantum machine learning systems, but it also provides an extensive overview of the state of research into Noise in Photonic Quantum Machine Learning, and identifies significant obstacles and future avenues of research for realizing feasible, scalable and effective Photonic Quantum Machine Learning Applications on the Photonic Quantum Machine Learning Platforms.

### 04 Apr 2026 v3

@@ 5.2 Advanced Characterisation Techniques
After characterizing the noise using the above techniques, we need to consider developing effective noise mitigation strategies to reduce the impact of noise on photonic quantum machine learning systems.
@@ 6.1 Hardware-Level Mitigation
Mitigation strategies at the hardware level focus on improving the quality of the physical components that comprise the photonic quantum system as a method for eliminating noise. One of the most effective methods to reduce noise is by incorporating quality optical components like optical waveguides with high-efficiency single-photon detectors and stable laser sources with low losses. By implementing low-loss waveguides, high-efficiency single-photon detectors, and stable laser sources, it becomes possible to effectively reduce the number of photons detected due to inefficiencies at the detector and lost to the environment due to extinction of the light [39]. Likewise, the incorporation of active stability techniques like feedback control systems will further enhance the stability of optical components, thereby eliminating phase noise [23].
@@ 6.2 Encoding-Based Mitigation
Encoding-based mitigation techniques are focused on creating a set of specific types of encoded agencies to help protect quantum information from degrading environmental disturbance or noise. One encoding method for quantum information is called Redundant Encoding, which is a way of spreading quantum information throughout multiple photons or modes of light within a physical composite quantum [9]. Redundant structured encoding spreads the quantum information out, providing more flexibility with regard to the loss of photons. The second form of encoding method is referred to as the use of Decoherence-Free Subspaces(DFS), which involves utilizing specific subspaces in the quantum region in order to encode the quantum information with immunity from a number of different forms of environmental noise, specifically forms of noise created through the combined dephasing of our universe [28]. Another encoding technique for encoding quantum information is the implementation of Quantum Error-Correcting Codes, which detect and correct errors that occur through environmental noise during quantum processes. Each of the above encoding techniques works to increase the resilience of quantum machine learning systems to a combination of different environmental noise sources.
@@ 6.3 Algorithmic Mitigation
Algorithmic mitigation techniques aim to build quantum machine learning algorithms that can stay resilient when faced with noise in photonic quantum systems. To that end, error mitigation methods are available; these are designed to help eliminate the effects of noise on the output of quantum algorithms and therefore do not have to perform complete error correction. Error mitigation methods include probabilistic error cancellation, zero-noise extrapolation, and symmetry verification [41]. Another approach used by algorithmic mitigation is to create noise-resilient algorithms that have been created specifically to handle certain noise types well. Noise-aware training is a technique that allows a user to train an algorithm to understand what type of noise is present during its learning phase. By combining these algorithmic mitigation techniques, we are able to improve quantum machine learning algorithm performance despite the presence of noise. Other than above approaches, Machine Learning-Based Noise Reduction techniques can also be used to reduce noise in photonic quantum systems. These techniques use machine learning models to identify and filter out noise from quantum data, thereby improving the quality of the quantum information being processed [43]. “Figure. 6” illustrates such a machine-learning-based quantum error mitigation approach.
Figure 6: Schematic illustration of machine learning-based quantum error mitigation where a trained model learns to map noisy expectation values obtained from a quantum circuit to noise-mitigated predictions through supervised training.
@@ 6.4 Hybrid Classical-Quantum Approaches
Hybrid classical-quantum approaches are strategies that merge classical and quantum computing technologies to reduce this interference. A hybrid classical-quantum approach can also be achieved through Classical Post-Processing, which uses classical algorithms for analysis and correction of the output data from quantum algorithms that have been degraded by noise as a result of their execution in a noisy environment [41]. On the other hand, variational hybrid algorithms can also be hybrid algorithms based on a combination of quantum computing and classical optimization techniques to develop adaptive algorithms that can continually adjust to the noise throughout their execution [40]. Hybrid classical-quantum approaches use a combination of the advantages of both classical and quantum computing to develop effective ways to reduce the effect of noise in the development of photonic quantum machine learning systems.
@@ 7 Recent Advances and Experimental Demonstrations
2016 was a seminal year for continuous-variable (CV) photonic quantum computing, as it was the year in which a 1 million-mode photonic cluster state was demonstrated, thus demonstrating the scale-up capabilities of CV photonic systems [44]. Using this very large-scale cluster state, a variety of quantum information processing tasks were also demonstrated (e.g., quantum teleportation and entanglement distribution).
Starting in 2021, research teams have developed integrated programmable photonic processors at multiple institutions (Xanadu, NIST, and MIT), thus enabling multiple quantum algorithms to be implemented on a single photonic chip, which enables reconfigurable and scalable photonic quantum computing [6]. In addition, the integrated programmable photonic processors also reduced system noise by decreasing both the number of optical components and improving the stability of the overall system. “Figure. 7” illustrates an overview of a processor.
Figure 7: System-level schematic of an integrated programmable photonic quantum processor, showing the interaction between a classical control interface, control electronics, photonic chip, optical components, and detection system, with feedback for system optimization.
@@
To continue developments toward 2024, a new superconducting-photonics hybrid chip was conceived, allowing for  99% loss resilience at photon levels and leading to the reduction of the leading type of noise in photonic systems [26]. These developments will allow for a much larger scale of practical deployment of photonic-based quantum computing.
@@ 8 Discussion
According to the above sections, we can discuss three main points, such as theoretical frameworks and noise modeling perspectives, near-term applications of PQML under noise, and long-term prospects and open challenges. These models are commonly expressed using quantum channels such as the Kraus representation introduced earlier.
@@ 8.1 Theoretical Frameworks and Noise Modeling Perspectives
To develop the most promising strategies to reduce or eliminate noise in photonic quantum systems, we must construct a theoretical foundation that accurately captures the types of noise found in these systems. Noise may have many different effects on the state of a quantum system; thus, there are several types of noise models that have been proposed to define how it impacts a quantum state, such as amplitude and phase damping and depolarizing channels [11]. By accurately modeling the various noise processes affecting photonic quantum systems, we may be able to predict how each noise type will act upon different types of quantum algorithms, as well as design robust quantum algorithms that are resilient to all types of noise. Furthermore, advanced techniques for modeling noise, including the development of stochastic noise models and non-Markovian noise models [10], provide the most accurate representation of the many types of noise that exist in the real world and will allow us to study the temporal correlations in noise, the memory effects associated with it, and the long-term behavior of quantum states in noisy environments.
@@ 8.2 Near-Term Applications of PQML under Noise
Advances in photonic quantum machine learning (PQML) will bring about a transformational impact in many areas, including but not limited to areas that are susceptible to noise. For example, due to its nature of being able to classify chemicals, photos, videos, etc. in real time, PQML could allow for more precise image classification than classical classification techniques [31]. A good example of this would be using a PQML algorithm for classifying chemicals found in samples of murky water, which cannot be done effectively by current methods of classification. As PQML continues to evolve, researchers will find even greater opportunities to utilize the incredible capabilities of PQML to their advantage, particularly due to the inherent robustness of PQML digital circuitry and room temperature operation.
@@ 8.3 Long-Term Prospects and Open Challenges
The future of photonic quantum machine learning (PQML) looks bright, but many current challenges are preventing this field from reaching its maximum potential. Among these challenges are the limitations on the scalability of quantum photon systems. Although considerable work has been completed in building integrated photonic circuits, creating systems that may undertake more qubits or continue functioning at a low noise level when built larger will be difficult to develop [6]. Additionally, integrating photonic quantum systems with other methods for executing quantum technologies such as superconducting qubit systems or an ion confined to a magnetic field presents not only opportunities to use the advantages of each unique quantum scheme but also creates additional issues associated with noise control and system coherence [5]. Therefore, addressing these challenges will require increased interdisciplinary cooperation among physicists, engineers, and computer scientists to advance the state-of-the-art of PQML further.
@@ 9 Conclusion
Our paper provided a comprehensive review of the state of research into quantum machine learning (QML) and photonic quantum computing (PQC) that focused on the noise limitation in the quantum machine learning systems. We have covered the general principles related to photonic quantum machine learning, including the various types of photonic quantum computing architectures, algorithms developed using quantum machine learning on photonic platforms, and various noise sources in photonic quantum systems and their classifications, including natural noise mechanisms and the characteristics associated with the photonic quantum systems. Also, we provided an overview of how the various forms of noise affect the quantum machine learning algorithm/algorithms and discussed the current research into algorithm-specific noise sensitivity and the design of noise-resilient algorithms.
Additionally, this review has included an overview of the various types of methods and current status of research into the various methods and current research into noise characterization, as well as how the various techniques, including hardware-level techniques to develop encoding-based, algorithmic, and hybrid classical-quantum methods, and experimental demonstrations work. Overall, this article highlights not only the challenges and achievements in managing the problems caused by noise within photonic quantum machine learning systems, but it also provides an extensive overview of the state of research into noise in photonic quantum machine learning and identifies significant obstacles and future avenues of research for realizing feasible, scalable, and effective photonic quantum machine learning applications on the photonic quantum machine learning platforms.